OpenStudy (anonymous):

Find the volume of the prism. A. 40.5 m3 B. 162 m3 C. 9 m3 D. 81 m3

OpenStudy (anonymous):



ganeshie8 (ganeshie8):

nice drawing :D

ganeshie8 (ganeshie8):

use this formula : volume = (base area) * height

ganeshie8 (ganeshie8):

base area = \(\frac{1}{2}b*h = \frac{1}{2}3*6\)